### v2

# When Fusion Helps and When It Breaks: View-Aligned Robustness in Same-Source Financial Imaging

Abstract
We study same-source multi-view learning and adversarial robustness for next-day direction prediction using two deterministic, window-aligned image views derived from the same time series: an OHLCV-rendered chart (ohlcv) and a technical-indicator matrix (indic). To control label ambiguity from near-zero moves, we use an ex-post minimum-movement threshold min_move ($\tau$) based on realized $|r_{t+1}|$, defining an offline benchmark on the subset $|r_{t+1}|\geq\tau$. Under leakage-resistant time-block splits with embargo, we compare early fusion (channel stacking) and dual-encoder late fusion with optional cross-branch consistency. We then evaluate pixel-space $\ell_{\infty}$ evasion attacks (FGSM/PGD) under view-constrained and joint threat models. We find that fusion is regime dependent: early fusion can suffer negative transfer under noisier settings, whereas late fusion is a more reliable default once labels stabilize. Robustness degrades sharply under tiny budgets with stable view-dependent vulnerabilities; late fusion often helps under view-constrained attacks, but joint perturbations remain challenging.
keywords:
multi-view learning , same-source , fusion , adversarial robustness , financial imaging , view-aligned evaluation
[…]
[inst1]organization=Xingchong (Xi’an) Technology Co., Ltd., addressline=R&D Department, city=Xi’an, state=Shaanxi, country=China
1 Introduction
More views do not automatically translate into better robustness. Financial time series prediction is notoriously sensitive to label ambiguity, non-stationarity, and evaluation leakage. In particular, next-day direction labels can be dominated by near-zero moves, where the notion of “up” versus “down” is weakly defined and easily overwhelmed by noise. At the same time, there is growing interest in vision-style representations of market data—candlestick charts, indicator heatmaps, and other renderings—because they enable the reuse of mature computer-vision architectures and data-augmentation pipelines. However, when multiple views or modalities are involved, practical comparisons are often confounded by cross-source alignment issues, inconsistent timestamps, or missingness patterns, making it difficult to attribute gains to fusion design rather than data integration.
This work studies same-source multi-view learning and adversarial robustness in a controlled setting. We construct two deterministic, window-aligned image views from the same underlying time series: (i) an OHLCV-rendered price/volume chart (ohlcv) and (ii) a technical-indicator matrix view (indic). Because both views are derived from identical timestamps and a shared window definition, the setup isolates representation and fusion effects from cross-source alignment confounds. We benchmark standard convolutional vision backbones and compare early fusion (channel stacking) against dual-encoder late fusion with a fusion head, optionally augmented with a cross-branch consistency regularizer.
A key methodological issue is that the unconditional next-day direction task can be nearly unlearnable when labels are dominated by micro-moves. To study learnability and robustness under controlled label-ambiguity regimes, we use an ex-post minimum-movement threshold min_move ($\tau$) defined by the realized next-day return magnitude $|r_{t+1}|$. This label-dependent filtering changes the task from unconditional direction prediction to direction prediction conditional on $|r_{t+1}|\geq\tau$; it is used solely for offline benchmarking and is not an inference-time decision rule. To avoid leakage and keep comparisons controlled, we apply filtering once prior to splitting and then run the embargoed time-block split on the filtered index set; details are in Appendices˜D and D.
We further evaluate adversarial robustness under an explicit view–channel threat model. Using pixel-space $\ell_{\infty}$ attacks on the rendered inputs (FGSM/PGD) [1, 2], we distinguish between view-constrained attacks that perturb exactly one view and joint attacks that perturb both views simultaneously. This protocol makes it possible to ask whether multi-view learning provides robustness through redundancy when only one view is corrupted, and whether such redundancy persists under coordinated perturbations. We also conduct operating-point sensitivity checks (e.g., at $\tau=0.008$) to assess which qualitative findings are stable and which are operating-point dependent (Appendix˜I).
Our main findings, summarized in Section˜6.2 and discussed in Section˜7, can be distilled as follows: (i) fusion behavior is strongly regime dependent—early fusion can exhibit negative transfer under noisier settings, while late fusion is a more reliable default once labels stabilize; (ii) naturally trained models degrade sharply under tiny pixel-space budgets, and robustness is markedly view dependent, revealing stable view asymmetries; (iii) late fusion often degrades more gracefully under view-constrained attacks, but joint perturbations remain challenging; and (iv) consistency regularization primarily improves cross-branch alignment, while its effects on clean performance and robustness are conditional and can shift with $\tau$.
Contributions
We make the following contributions:
  • 1.
    Controlled same-source multi-view benchmark. We define a window-aligned dual-view imaging protocol for SGE gold spot data (2005–2025), enabling clean attribution of fusion effects without cross-source alignment confounds.
  • 2.
    Operating-point-aware evaluation. We use min_move to construct controlled label-ambiguity regimes for offline benchmarking, paired with leakage-resistant time-block splits with embargo to reduce temporal leakage (Sections˜5.4 and D).
  • 3.
    View–channel adversarial robustness analysis. We evaluate FGSM/PGD under view-constrained and joint threat models, characterize sharp degradation under tiny budgets, and document stable view-dependent vulnerabilities and the conditional benefits of late fusion (Section˜6.2).
Finally, we note the scope of this study: our baselines focus on convolutional vision backbones (Lite-CNN and ResNet18-P) and pixel-space attacks on rendered inputs. Extending the benchmark to Transformer-based architectures and broader threat models remains an important direction for future work.
Organization
After reviewing related work, we introduce the same-source multi-view setup, task, and evaluation protocol, and then specify the threat model and view-aligned robustness evaluation in Section˜4. Clean baselines and fusion behavior are reported in Section˜6.1, followed by adversarial robustness results in Section˜6.2. We discuss implications and limitations in Section˜7 and conclude in Section˜8. Additional implementation details and extended analyses—including the $2/255$ stress test, operating-point sensitivity at $\tau=0.008$, and full branch diagnostics—are provided in Appendices˜J, I and H.
Code availability
Code to reproduce the experiments will be released upon acceptance at: https://github.com/mrui166/same-source-financial-imaging-robustness.
2 Related Work
2.1 Financial Image Representations and Multi-View Fusion for Market Prediction
A common line of work in market prediction converts financial time series into 2D visual representations so that convolutional models can exploit local spatial patterns and cross-variable structures. Typical examples include candlestick-/OHLCV-style charts and chart-pattern based encodings, as well as indicator-driven matrices/heatmaps constructed from technical signals. Beyond ad-hoc chart rendering, a more principled “time-series-to-image” family encodes temporal structure into images such as Gramian Angular Fields and Markov Transition Fields, enabling multi-channel image learning from a single sequence [3]. In financial settings, such imaging ideas have been instantiated in different ways, including technical-indicator image grids for trading [4], bar-/chart-window image learners [4, 5], and candlestick-pattern imaging for CNN-based recognition [6]. Related work also considers incorporating cross-asset market structure via graph-enhanced predictors, e.g., a graph convolutional feature based CNN for stock trend prediction [7]. Recent surveys often group multi-source fusion methods for financial market prediction by when fusion is applied—pre-model (early/data), mid-model, or post-model (late/decision)—and note that predictive performance can be sensitive to this choice [8]. Related strands also explore learning from different representations of the same underlying data, sometimes fusing them (e.g., feature-level fusion in hybrid CNN–RNN pipelines) [9] or incorporating pattern-based representations such as candlesticks and sequence similarity [10].
Within this landscape, our two-view construction follows the multi-view intuition that complementary representations can provide partially redundant evidence. We therefore treat complementarity as an empirical question and explicitly analyze regime dependence under label-stabilization settings (via a minimum-movement threshold). In particular, we adopt an OHLCV-rendered view that preserves price-action/volume cues and an indicator-matrix view that emphasizes trend/momentum signals; our same-source two-view design is inspired by prior multi-source image-integration approaches in financial forecasting [11], while our focus is on the same-source case to remove cross-source alignment confounds. We focus on a same-source setting where both views are deterministically derived from an aligned rolling window of a single market time series (SGE gold spot), which isolates the role of representation and fusion from confounding cross-source alignment issues.
A key practical takeaway from prior fusion work is that multi-view learning is not automatically beneficial: gains depend on whether the views are complementary and on careful representation alignment and fusion design [12, 13]. More broadly, financial forecasting pipelines are also prone to subtle time leakage (e.g., through rolling normalization, window overlap, or hyperparameter search) and selection effects from repeated backtests; protocol designs such as purged CV/embargo and diagnostics such as PBO/DSR have been advocated to diagnose and quantify these risks [14, 15, 16]. Related concerns about leakage undermining benchmark validity have also been discussed in other areas (e.g., efforts to mitigate memorization-driven leakage in LLM QA benchmarks) [17], which serves as a reminder that evaluation reliability depends on stating leakage-related assumptions explicitly. Motivated by this, our study emphasizes a controlled comparison between early fusion (channel stacking) and late fusion (dual encoders with a fusion head), and interprets gains in an evaluation-reliability–oriented context of leakage-resistant time splits (Section˜5.4) and label-stabilization regimes (Section˜5.3). Many closely related financial-imaging papers report improvements under relatively controlled evaluation, whereas we center the analysis on evaluation protocol choices (e.g., leakage control, operating-point selection) and fusion paradigms as first-class factors that shape reliability.
2.2 Adversarial Robustness and Threat Models for Multi-View Inputs
Adversarial robustness has been extensively studied in vision, where small $\ell_{\infty}$-bounded perturbations can induce systematic misclassification [18, 1]. Canonical first-order attacks include FGSM [1] and multi-step PGD [2], which have become standard stress tests for probing worst-case sensitivity under first-order adversaries. Beyond generic vision benchmarks, a smaller but growing line of work considers adversarial sensitivity in financial vision models built on chart-like representations. For example, Chen et al. study adversarial robustness of a deep convolutional candlestick learner built on Gramian Angular Fields (GAF) encodings and propose perturbation sampling/adversarial training to improve stability under structured perturbations of candlestick inputs [19]. These studies are complementary to ours: they focus on candlestick-pattern classification and representation-specific attacks/defenses, whereas we evaluate next-day direction prediction under leakage-resistant time splits and explicitly compare view-constrained vs. joint threat models for same-source two-view fusion. Beyond single-input settings, robustness questions also arise for multi-branch and multimodal models, where the threat model can be defined per input stream (single-source/view-constrained attacks), and where fusion architecture can modulate robustness and failure modes [20]. In our setting, we evaluate an inference-time, white-box, untargeted FGSM/PGD threat model instantiated as raw pixel-space $\ell_{\infty}$-bounded perturbations on rendered views, with optional standardized-space execution via budget conversion that preserves the underlying raw-space constraint. We emphasize that this pixel-space threat model is a representation-level stress test of worst-case sensitivity for the chosen rendering and fusion design, rather than a claim about a realistic financial adversary operating in the underlying market or data-generation process.
Time-series settings introduce additional subtleties: perturbations can be defined either in the raw temporal domain or on downstream representations (e.g., rendered images), and the choice determines what “small” means and which invariances are being tested. Prior work has instantiated adversarial attacks directly on time-series classifiers and shown that standard gradient-based methods can transfer to temporal signals [21]. In financial pipelines, where inputs often undergo deterministic preprocessing (windowing, indicator computation, normalization), robustness evaluation therefore benefits from threat models that explicitly account for where perturbations are applied and how multi-view inputs are aligned.
Our work aligns with this perspective by treating pixel-space $\ell_{\infty}$ perturbations on rendered views as a representation-level stress test, and by formalizing the multi-view attack surface through view–channel mapping. This enables (i) view-constrained perturbations that target exactly one view in two-view models and (ii) joint perturbations that attack both views, while maintaining a view-aligned comparison to single-view baselines. Such protocols help disentangle robustness contributions from (a) the chosen representation, (b) early vs. late fusion, and (c) auxiliary alignment objectives (e.g., cross-view consistency), while reducing confounds from data-splitting artifacts or cross-source misalignment.
3 Problem Setup and Notation
3.1 Same-Source Data (SGE Gold Spot)
Source. Shanghai Gold Exchange (SGE), spot gold; data retrieved via Tushare [22].
Time range. 2005-09-26–2025-08-28 (all dates reported in UTC).
Granularity. Trading-day level observations (non-trading days excluded; trading days follow the SGE calendar).
Fields. Price and trading activity information (OHLC and trading volume).
Preprocessing. We remove invalid records, sort the data chronologically, and construct rolling lookback windows. Both image views used in our experiments are generated from the same window and aligned to the same window end date.
3.2 Multi-View Construction
Motivated by the multi-view formulation in MODII [11], we construct two complementary 2D image views from the same market-data window: (i) an OHLCV-based price/volume view (ohlcv) and (ii) a technical-indicator view (indic). Both views are generated from the same reference day and aligned to the same 15-day lookback window, enabling controlled comparisons across single-view and two-view settings. Our construction follows a MODII-style pipeline [11] and is conceptually related to prior time-series-to-image transformations for financial prediction [4]. Rendering, scaling, and ordering details are provided in B.
OHLCV view (ohlcv)
We render OHLC bars with a volume subpanel to capture recent level, range, and trading activity in a compact visual representation.
Indicator view (indic)
We compute a fixed set of $N=15$ standard indicators over the same lookback window and convert the resulting indicator-by-time matrix into a grayscale image; the indicator set and parameters are fixed a priori (B).
Figure 1: Examples of ohlcv and indic constructed from a fixed lookback window of $L_{\mathrm{ohlcv}}=L_{\mathrm{indic}}=15$ trading days.
3.3 View-to-Channel Mapping
For two-view models (both), we concatenate the two grayscale images along the channel dimension,
$\mathbf{x}=[v_{\texttt{ohlcv}};v_{\texttt{indic}}]\in\mathbb{R}^{2\times 256\times 256},$ (1)
[…]
3.4 Task and Evaluation Metrics
Prediction target
For each trading day $t$, we form an input sample by extracting a lookback window of the past $L$ trading days, $\{t-L+1,\dots,t\}$, and predict the next-day directional movement at $(t+1)$. In all experiments, we set $L=15$. We define the next-day simple return as
$r_{t+1}=\frac{close_{t+1}-close_{t}}{close_{t}},\qquad y_{t}=\mathbb{I}\!\left(r_{t+1}>0\right),$ (2)
where $close_{t}$ denotes the closing price on day $t$ and $\mathbb{I}(\cdot)$ is the indicator function. Thus, the task is a binary classification problem with labels $y_{t}\in\{0,1\}$. When using the ex-post movement threshold min_move ($\tau$), we report performance on the conditional subset $|r_{t+1}|\geq\tau$; see D for the filtering/split protocol.
Matthews Correlation Coefficient (MCC)
We use MCC as the primary metric because it summarizes the full confusion matrix and remains informative under class imbalance [23]. MCC can be interpreted as a correlation between predicted and true labels, taking values in [1,1][-1,1].
$\mathrm{MCC}=\frac{TP\cdot TN-FP\cdot FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}.$ (3)
[…]
4 Threat Model and Evaluation Protocol
We evaluate test-time, white-box evasion robustness on the rendered image inputs using standard pixel-space $\ell_{\infty}$ perturbations and first-order attacks [18, 1, 2]. Our goal is a view-aligned stress test: (i) perturb exactly one view for two-view models (view-constrained) or (ii) perturb both views jointly, while keeping comparisons to single-view baselines well defined via the view–channel mapping (Section˜3.3).
4.1 Threat Model
$\ell_{\infty}$-bounded perturbations in raw pixel space
Let $\mathbf{x}_{\mathrm{raw}}\in[0,1]^{C\times 256\times 256}$ denote the raw input (with $C=1$ for single-view and $C=2$ for two-view). An adversary produces $\mathbf{x}^{\prime}_{\mathrm{raw}}=\mathbf{x}_{\mathrm{raw}}+\boldsymbol{\delta}$ subject to [18, 1, 2]
$\|\boldsymbol{\delta}\|_{\infty}\leq\epsilon_{\mathrm{adv}},\qquad\|\boldsymbol{\delta}\|_{\infty}\triangleq\max_{i}|\delta_{i}|.$ (4)
Budgets
We evaluate a small-budget grid
$E=\left\{0,\ \frac{0.25}{255},\ \frac{0.5}{255},\ \frac{0.75}{255},\ \frac{1}{255}\right\},$ (5)
where $\epsilon_{\mathrm{adv}}=0$ corresponds to clean evaluation.
Attacks
We use FGSM and multi-step PGD as untargeted cross-entropy attacks [1, 2]. All attack hyperparameters (steps, step size, projection, and clipping) are reported in F.
Standardization and valid-range handling
The threat model is always defined in raw $[0,1]$ space. When models consume standardized inputs, we execute attacks in the standardized space using budget conversion that preserves the underlying raw-space constraint, and clip to the corresponding valid range (A).
4.2 Threat Scenarios and View-Aligned Protocol
Using the view–channel mapping (Section˜3.3), we define two threat scenarios.
View-constrained perturbations (single-view corruption)
For two-view inputs, the adversary perturbs exactly one view/channel [20]:
$\boldsymbol{\delta}=(\delta_{\texttt{ch0}},\mathbf{0})\ \text{(attack {ohlcv})}\quad\text{or}\quad\boldsymbol{\delta}=(\mathbf{0},\delta_{\texttt{ch1}})\ \text{(attack {indic})}.$ (6)
Joint perturbations (coordinated corruption)
For two-view inputs, the adversary perturbs both channels:
$\boldsymbol{\delta}=(\delta_{\texttt{ch0}},\delta_{\texttt{ch1}}).$ (7)
View-aligned comparison to single-view baselines
Single-view models always place the provided view in ch0; therefore, we align comparisons by the attacked view: ohlcv (single-view ohlcv attacked on ch0) vs. (both attacked on ch0), and indic (single-view indic attacked on ch0) vs. (both attacked on ch1). Joint perturbations apply only to two-view models.
4.3 Evaluation and Summary Metrics
[…]
$\Delta(\epsilon_{\mathrm{adv}})=\mathrm{MCC}(\epsilon_{\mathrm{adv}})-\mathrm{MCC}(0),\qquad\Delta_{\mathrm{worst}}=\min_{\epsilon_{\mathrm{adv}}\in E,\ \epsilon_{\mathrm{adv}}>0}\Delta(\epsilon_{\mathrm{adv}}).$ (8)
Late-fusion diagnostics (analysis only)
For late-fusion two-view models, we additionally report branch-level diagnostics on the clean test set to aid interpretation; see H.
5 A Stable, Leakage-Resistant Training Baseline
We consider a set of learning and non-learning baselines to evaluate our image representations under a unified, leakage-resistant evaluation protocol. All models perform binary next-day direction prediction and output logits/probabilities. Inputs are either a single view (indic or ohlcv) or a two-view input (both); for early-fusion models, the two resized views are stacked channel-wise (Section˜3.3). Unless otherwise stated, trainable baselines share the same optimization and checkpoint-selection protocol (E).
5.1 Model Architectures
Compact CNN baseline (Lite-CNN)
We use a lightweight convolutional network adapted from the MODII framework [11]. In both mode, we use early fusion by channel-wise stacking of the two views; in single-view mode, the model consumes one channel.
Pretrained ResNet-18 baseline (ResNet18-P)
We employ a ResNet-18 backbone initialized with ImageNet pretrained weights [24, 25]. We adapt only the first convolution to match the required number of input channels; the rest of the backbone is unchanged. Two-view both uses early fusion (stacked channels) as above.
Dual-view consistency variants (*-cons)
For two-view inputs, we additionally evaluate late-fusion dual-branch variants for both backbones, denoted Lite-CNN-Cons and ResNet18-P-Cons. These models produce view-specific logits and a fused-logit output; training uses the objective in Section˜5.2. We report $\lambda\in\{0,0.5,1\}$ (with $\lambda=0$ reducing to fused-logit cross-entropy only) and fix $T=2.0$ unless stated otherwise.
Logistic Regression baseline (LogReg)
We apply logistic regression on flattened image features to measure how far linear decision boundaries on the constructed representations can go.
Majority-class baseline (Majority)
We include a non-learning baseline that always predicts the most frequent class in the training split.
For reproducibility, implementation details of backbones and fusion heads are provided in C, including a concise implementation-level summary in Table˜4.
5.2 Cross-view Consistency Regularization
Motivation
In the two-view setting (both), each sample provides two aligned views of the same market state and shares the same label. We encourage the two view-specific predictors to produce coherent predictions via an explicit consistency regularizer.
Dual-branch setup
Let the two branches produce logits $z_{a}$ and $z_{b}$ from the two views, and let a fusion head produce fused logits $z_{f}$ from concatenated branch features. The fused logits $z_{f}$ define the supervised objective; $z_{a}$ and $z_{b}$ define the consistency term.
Temperature-scaled symmetric KL
We define temperature-scaled predictive distributions [26, 27]
$p_{a}=\mathrm{softmax}(z_{a}/T),\qquad p_{b}=\mathrm{softmax}(z_{b}/T),$ (9)
and use the symmetric KL divergence
$\mathcal{L}_{\mathrm{cons}}(z_{a},z_{b})=\frac{T^{2}}{2}\Big[\mathrm{KL}(p_{a}\,\|\,p_{b})+\mathrm{KL}(p_{b}\,\|\,p_{a})\Big].$ (10)
Training objective
We optimize
$\mathcal{L}=\mathcal{L}_{\mathrm{CE}}(z_{f},y)+\lambda\,\mathcal{L}_{\mathrm{cons}}(z_{a},z_{b}),$ (11)
where $\lambda\geq 0$ controls the regularization strength. Branch-level diagnostics used for interpretation are reported in H.
5.3 Stabilizing Labels with a Minimum-Movement Threshold
Next-day direction labels can be ambiguous when the realized price change is extremely small. To study how label ambiguity affects fusion and robustness, we optionally define an ex-post evaluation subset using a minimum-movement threshold $\tau$: for each sample ending at day $t$, we compute $r_{t+1}$ (Eq. 2) and keep the sample only if $|r_{t+1}|\geq\tau$. We evaluate $\tau\in\{0,\,0.002,\,0.004,\,0.006,\,0.008,\,0.010\}$ (denoted min_move).
Protocol note
Because inclusion depends on the realized $|r_{t+1}|$, min_move defines a conditional offline benchmark rather than an inference-time decision rule. To keep comparisons controlled across split variants, filtering is applied before splitting so that all variants operate on the same subset definition (D).
5.4 Leakage-Resistant Time Split with Embargo
Because inputs are built from overlapping sliding windows, random splitting can introduce look-ahead leakage. We therefore adopt a leakage-resistant time-block split with embargo for all models [14]. Trading days are grouped into consecutive blocks of size $B$ (block_size); earlier blocks form train, then validation, and the latest blocks form test (default 70/15/15 by blocks).
To mitigate boundary leakage across split boundaries, we apply an embargo of at least $D_{\mathrm{emb}}$ trading days (embargo_days[14, 28]. Implementation details and the sample-size impact under our default configuration (block_size=30, embargo_days=20) are reported in D (Table˜5).
5.5 Implementation Details
Training and selection details (optimizer, early stopping, and reproducibility controls) are provided in E. Input standardization and leakage-avoiding normalization are detailed in A, and adversarial-attack hyperparameters are summarized in F.
6 Results
6.1 Clean Baselines
[…]
Setup and evaluation metric
We compare Majority, LogReg, and two CNN backbones (Lite-CNN and ResNet18-P) under three input modes: ohlcv, indic, and dual-view both. For both, vanilla CNN baselines use early fusion (channel-wise stacking), whereas *-cons variants use a dual-branch late-fusion head and optionally add cross-view consistency regularization weighted by $\lambda$. We use MCC as the primary decision metric since it accounts for all TP/TN/FP/FN entries and penalizes degenerate single-direction predictors.
Minimum-movement filtering determines the effective difficulty
As shown in Figure˜2, performance is near chance without filtering or under weak filtering ($\tau\in\{0,0.002,0.004\}$), but improves sharply once $\tau$ reaches the mid range (notably $\tau\in\{0.006,0.008\}$), consistent with reduced label ambiguity on the ex-post filtered subsets. Aggregating across all deep configurations (Lite-CNN/ResNet18-P and *-cons, all input modes), the mean MCC rises from 0.002-0.002 at $\tau=0$ to $0.082$ at $\tau=0.006$ and peaks at $0.092$ at $\tau=0.008$, before dropping to $0.030$ at $\tau=0.010$. The drop at $\tau=0.010$ aligns with the reduced effective sample size and increased variance (see the sample-count/embargo summary in Table˜5, in Appendix˜D). We stress that $\tau$ defines offline evaluation regimes ex post, not an inference-time filtering rule; the filtering/splitting protocol is specified in Appendix˜D.
Fusion behavior once labels stabilize
Once $\tau\geq 0.006$, dual-view information becomes exploitable and both turns consistently positive for Lite-CNN family; however, under weak filtering (e.g., $\tau=0.004$) early fusion can exhibit negative transfer relative to the stronger single-view option (Table˜1). Across the two-view setting, late fusion is the dominant architectural gain over early fusion (representative numbers in Table˜1), while consistency regularization yields modest, $\tau$-dependent and non-monotonic effects.
Backbone differences
ResNet18-P exhibits a noisier regime: dual-view early fusion is generally weak relative to single-view modes at the same $\tau$, and late fusion partially recovers performance but with higher variance, especially under stricter filtering (see Figure˜2 and representative values in Table˜1).
LogReg as a diagnostic
LogReg can be competitive at certain thresholds but is non-monotonic across $\tau$ and input modes; we treat it primarily as a diagnostic indicator of partial linear separability after label stabilization.
Takeaways
Across backbones, (i) label stabilization via min_move is a prerequisite for learning in our setting; (ii) in the dual-view setting, late fusion is generally more reliable than early fusion on clean evaluation once labels stabilize; and (iii) consistency regularization can provide modest gains in some mid-noise regimes but is not reliably monotonic and requires careful tuning jointly with $\tau$. For full numerical commentary (without changing any conclusions or values), see Appendix˜G.
[…]
Model (input) $\boldsymbol{\tau=0.004}$ $\boldsymbol{\tau=0.006}$ $\boldsymbol{\tau=0.008}$ Lite-CNN (both, early fusion) $020004\pm 0.016396 0.0$$088767\pm 0.042662$ $\mathbf{0.107206\pm 0.046632}$ Lite-CNN (indic) $\mathbf{0.107292\pm 0.022477}$ $0.079303\pm 0.032678$ $0.100279\pm 0.048322$ Lite-CNN (ohlcv) $0.003236\pm 0.023789$ $109095\pm 0.039023 0.0$$090707\pm 0.028893$ Lite-CNN-Cons (both, $\lambda=0$) $0.023262\pm 0.034602$ $0.127459\pm 0.035946$ $0.100430\pm 0.034439$ Lite-CNN-Cons (both, $\lambda=0.5$) $0.015634\pm 0.052001$ $\mathbf{0.129607\pm 0.026449}$ $0.089973\pm 0.067209$ Lite-CNN-Cons (both, $\lambda=1$) $0.014013\pm 0.034309$ $0.120755\pm 0.029468$ $0.103269\pm 0.031387$ ResNet18-P (both, early fusion) $012168\pm 0.030674 0.0$$17593\pm 0.044072 0.0$$042873\pm 0.061370$ ResNet18-P (indic) $055638\pm 0.047824 0.0$$082039\pm 0.039988$ $0.108023\pm 0.040145$ ResNet18-P (ohlcv) $005656\pm 0.059240 0.0$$018580\pm 0.054761$ $0.109101\pm 0.072830$ ResNet18-P-Cons (both, $\lambda=0.5$) $003653\pm 0.037078 0.0$$67112\pm 0.031388 0.0$$083423\pm 0.052584$
Choice of $\tau$ for robustness evaluation
All subsequent adversarial evaluations use $\tau=0.006$ as the default operating point. At smaller thresholds ($\tau\leq 0.004$), most configurations remain close to chance in MCC, making robustness comparisons less informative. At larger thresholds (e.g., $\tau=0.010$), the reduced effective sample size can amplify variance across random seeds. As a sensitivity check, we repeat the key robustness evaluations at $\tau=0.008$ and observe the same qualitative conclusions (see Appendix˜I). Thus, $\tau=0.006$ offers a practical balance between label stabilization and sufficient data for reliable robustness measurement.
6.2 Adversarial Robustness
Protocol and summary metrics
[…]
Summary of findings
At the default operating point $\tau=0.006$, we observe (i) rapid degradation under tiny pixel-space budgets, (ii) pronounced view dependence that is backbone- and fusion-dependent, and (iii) a robustness advantage of late fusion under view-constrained attacks, while robustness under joint (two-view) attacks remains challenging. The sensitivity study at $\tau=0.008$ in Appendix˜I confirms the structural trends in (i) and (iii), but indicates that finer-grained hyperparameter conclusions (e.g., the best $\lambda$ within late fusion) can shift with the operating point. In contrast, the direction of view asymmetry is primarily model- and fusion-dependent: ResNet18-P is consistently most vulnerable to indic-only perturbations, whereas Lite-CNN-Cons in late fusion is consistently more vulnerable to ohlcv-only perturbations under view-constrained attacks (Tables˜2 and 11).
[…]
Overall vulnerability under tiny $\ell_{\infty}$ budgets
Across architectures, naturally trained models can degrade substantially under pixel-space perturbations. Even at $\epsilon_{\mathrm{adv}}=0.25/255$ (a quarter of an 8-bit intensity level), PGD often causes a pronounced MCC drop, and performance can become near-zero or negative as the budget increases toward $1/255$. This confirms that pixel-space $\ell_{\infty}$ robustness is not obtained “for free” in our financial-image setting.
View sensitivity and view-aligned evaluation
Using the view–channel mapping in Section˜3.3, we compare (i) single-view models attacked on their sole channel and (ii) two-view models attacked on the corresponding view channel (Section˜4.2). The results reveal strong view dependence, but its direction is backbone- and fusion-dependent. For ResNet18-P, perturbing the indicator view is consistently the most destructive at both $\tau=0.006$ and $\tau=0.008$, and very small budgets can already drive MCC close to the lower bound, suggesting near-saturated failure rather than a gradual attenuation of predictive signal. For Lite-CNN variants, view asymmetry is milder and depends on the fusion pathway; notably, for Lite-CNN-Cons with late fusion, ohlcv-only perturbations are consistently the most destructive single-view attacks at both $\tau$ values (Tables˜2 and 11).
Late fusion improves robustness under view-constrained attacks
A consistent trend in the two-view setting is that dual-branch late fusion (*-cons) degrades more gracefully than early fusion (channel stacking) under view-constrained attacks. When only one view is perturbed, the unperturbed branch can still provide usable evidence, which mitigates worst-case collapse and yields a flatter robustness curve over $E$. This advantage is most pronounced in the Lite-CNN family under single-view attacks (and persists at $\tau=0.008$; Appendix˜I). For ResNet18-P, late fusion improves robustness against OHLCV-only perturbations but does not prevent saturated failures under indicator-only attacks.
Joint perturbations remain challenging
When the adversary perturbs both channels jointly, performance deteriorates more rapidly than in view-constrained settings, even for late fusion. This indicates that redundancy across views provides meaningful protection against single-view corruption, but does not fully defend against worst-case joint perturbations.
Consistency regularization: alignment, not a universal robustness knob
Within late-fusion models, the consistency weight $\lambda$ primarily enforces cross-branch logit alignment, while its impact on fused-head robustness is secondary and can be non-monotonic across operating points. We therefore treat $\lambda$ as a tunable refinement on top of the late-fusion design rather than a universally optimal setting; detailed branch diagnostics supporting this interpretation are provided in Appendix˜H.
Table 2: PGD robustness summary at $\tau=0.006$ (min_move). We report MCC at two representative budgets ($0.25/255$ and $1/255$) and the worst-case drop $\Delta_{\mathrm{worst}}$ over nonzero budgets $\{0.25,0.5,0.75,1\}/255$ (Eq. (8)). As an additional stress test, results at $\epsilon_{\mathrm{adv}}=2/255$ are reported in Appendix˜J.
Model Input Attack (PGD) MCC@0 MCC@0.25/255 MCC@1/255 $\boldsymbol{\Delta_{\mathrm{worst}}}$ Lite-CNN ohlcv on ohlcv 0.109 0.054 -0.088 -0.197 Lite-CNN indic on indic 0.079 0.024 -0.186 -0.265 Lite-CNN both (early) on ohlcv (ch0) 0.089 0.030 -0.104 -0.193 Lite-CNN both (early) on indic (ch1) 0.089 0.033 -0.094 -0.183 Lite-CNN both (early) on both 0.089 -0.014 -0.317 -0.406 Lite-CNN-Cons ($\lambda{=}1$) both (late) on ohlcv (ch0) 0.121 0.081 -0.029 -0.150 Lite-CNN-Cons ($\lambda{=}1$) both (late) on indic (ch1) 0.121 0.113 0.084 -0.037 Lite-CNN-Cons ($\lambda{=}1$) both (late) on both 0.121 0.071 -0.072 -0.192 ResNet18-P ohlcv on ohlcv 0.019 -0.443 -0.952 -0.971 ResNet18-P indic on indic 0.082 -0.999 -1.000 -1.082 ResNet18-P-Cons ($\lambda{=}1$) both (late) on ohlcv (ch0) 0.095 -0.211 -0.774 -0.870 ResNet18-P-Cons ($\lambda{=}1$) both (late) on indic (ch1) 0.095 -0.959 -1.000 -1.095 ResNet18-P-Cons ($\lambda{=}1$) both (late) on both 0.095 -0.983 -1.000 -1.095
Takeaways
Across all settings, pixel-space $\ell_{\infty}$ perturbations can severely degrade next-day direction prediction even at extremely small budgets. Robustness is strongly view-dependent and architecture-specific (ResNet18-P is particularly fragile to indic-only attacks). In the two-view regime, dual-branch late fusion generally improves robustness under view-constrained attacks, supporting the intuition that an unperturbed branch can provide a partial evidence path. However, joint perturbations remain challenging and can still induce large performance drops. Finally, consistency regularization mainly controls cross-branch alignment; its effect on robustness is secondary and can depend on the operating point (Appendix˜I).
PGD stress test at $\epsilon_{\mathrm{adv}}=2/255$ (not in the main grid)
We extend the perturbation budget to $\epsilon_{\mathrm{adv}}=2/255$ for late-fusion two-view Lite-CNN-Cons models and find that they remain non-trivially robust under the view-constrained indic-only attack (ch1), while ohlcv-only and joint attacks are substantially more destructive; full results at both $\tau=0.006$ and $\tau=0.008$ are reported in Appendix˜J.
7 Discussion
Operating point, noise regime, and fusion
All findings should be interpreted relative to the operating point induced by min_move ($\tau$) in Section˜6.1. Increasing $\tau$ reduces direction-label ambiguity by filtering near-zero moves but also reduces effective sample size and can increase variance. In noisy regimes (weak filtering), early fusion by channel stacking can underperform the stronger single-view baseline, consistent with negative transfer; once labels stabilize (around $\tau\geq 0.006$ in our setup), using both views becomes more reliably beneficial and dual-branch late fusion is a robust default. Sensitivity at $\tau=0.008$ in Appendix˜I preserves the qualitative patterns, while showing that some within-family hyperparameter preferences can shift with the operating point.
Robustness under tiny budgets: redundancy helps only under view-constrained attacks
Across backbones and operating points, naturally trained models can degrade sharply under small pixel-space $\ell_{\infty}$ budgets (Sections˜6.2, 2 and 11). Late fusion often degrades more gracefully under view-constrained attacks, consistent with preserving an unperturbed evidence path when only one view is perturbed. This mitigation is limited: view-specific failure modes can still saturate quickly, and joint perturbations remain challenging for both early and late fusion, including in the $2/255$ stress test (Appendix˜J).
Stable view asymmetry as a diagnostic
Vulnerability is strongly view dependent. ResNet18-P is consistently most fragile to indicator-view perturbations, while Lite-CNN-Cons under late fusion is more sensitive to ohlcv-only perturbations under view-constrained attacks. This suggests that robustness depends on view rendering and backbone inductive bias, and the direction of view asymmetry provides a useful diagnostic for interpreting view-aligned robustness results and identifying view-specific failure modes.
Consistency regularization and limitations
Branch diagnostics in Appendix˜H show that increasing $\lambda$ improves cross-branch agreement, but robustness effects are conditional: the clean-optimal $\lambda$ is backbone-dependent at $\tau=0.006$ (Section˜6.1), and robustness rankings among $\lambda\in\{0,0.5,1\}$ can change with $\tau$ (Appendix˜I). Our study evaluates pixel-space attacks on rendered images (a representation-level stress test), uses a label-dependent subset definition for offline benchmarking (Appendix˜D), and benchmarks only convolutional backbones; broader model families and threat models remain valuable extensions.
8 Conclusion
We studied same-source multi-view learning and adversarial robustness for next-day direction prediction using two deterministic, window-aligned image views (ohlcv and indic), enabling explicit view–channel threat modeling. Under leakage-resistant time-block splits with embargo, learnability and robustness depend strongly on the label-ambiguity regime induced by near-zero moves. We therefore use an ex-post minimum-movement filter (min_move) to define conditional evaluation subsets, revealing a non-monotonic data–noise trade-off: moderate $\tau$ improves learnability on the remaining subset, while overly strict filtering reduces sample size and increases variance. Importantly, min_move is a label-dependent subset definition for offline benchmarking and is not an inference-time decision rule (Appendix˜D).
In stabilized regimes (default $\tau=0.006$), early fusion can suffer negative transfer under noisier settings, whereas dual-branch late fusion is a more reliable default and provides the dominant clean-performance gains (Section˜6.1). From a robustness perspective, naturally trained models can be highly vulnerable to pixel-space $\ell_{\infty}$ perturbations even at tiny budgets (Section˜6.2). Late fusion often degrades more gracefully under view-constrained attacks, but joint perturbations remain challenging, indicating that cross-view redundancy alone does not guarantee worst-case robustness (Appendix˜J). Sensitivity at $\tau=0.008$ preserves the qualitative conclusions while highlighting operating-point dependence in within-family hyperparameters (Appendix˜I). Overall, reliable robustness assessment in same-source financial imaging benefits from leakage-resistant temporal evaluation, explicit reporting of operating points and label-dependent subset definitions, and threat models that distinguish view-constrained from joint multi-view attacks.
Declaration of competing interests
The author declares no competing interests.
Appendix A Implementation Details for Input Parameterization
Table˜3 summarizes the input parameterization and normalization settings.
Raw input
Each view is stored as an 8-bit grayscale image and loaded with ToTensor(), yielding a raw input $\mathbf{x}_{\mathrm{raw}}\in[0,1]^{C\times 256\times 256}$, where $C=1$ for single-view and $C=2$ for two-view models.
Optional per-channel standardization
Optionally, we apply per-channel standardization
$\mathbf{x}_{\mathrm{norm}}=(\mathbf{x}_{\mathrm{raw}}-\boldsymbol{\mu})\oslash\boldsymbol{\sigma},$ (12)
where $(\boldsymbol{\mu},\boldsymbol{\sigma})$ are estimated by randomly sampling $\min(n_{\text{norm}},|{\rm train}|)$ images from the training split (default $n_{\text{norm}}=512$) without replacement, computing per-channel statistics over $N\times H\times W$, and clamping $\sigma$ to be at least $10^{-6}$. When use_norm is enabled, the model consumes $\mathbf{x}_{\mathrm{norm}}$; otherwise, it consumes $\mathbf{x}_{\mathrm{raw}}$.
Valid-range clipping
Clipping is performed in the space where the attack is executed: raw inputs are clipped to $[0,1]$; standardized inputs are clipped to $\big[(0-\boldsymbol{\mu})\oslash\boldsymbol{\sigma},\ (1-\boldsymbol{\mu})\oslash\boldsymbol{\sigma}\big]$.
Threat model defined in raw space
The threat model is always defined on $\mathbf{x}_{\mathrm{raw}}$ with an $\ell_{\infty}$ constraint $\|\boldsymbol{\delta}\|_{\infty}\leq\epsilon_{\mathrm{adv}}$, where $\|\boldsymbol{\delta}\|_{\infty}=\max_{i}|\delta_{i}|$. When the model consumes standardized inputs, we enforce the same underlying raw-space budget by converting
$\epsilon_{\mathrm{norm}}=\epsilon_{\mathrm{adv}}\oslash\boldsymbol{\sigma},\qquad\alpha_{\mathrm{norm}}=\alpha\oslash\boldsymbol{\sigma},$ (13)
and then performing clipping/projection in the standardized space. This preserves the raw-space bound after mapping back.
Table 3: Input parameterization and normalization details used in our implementation.
Item Setting Image storage 8-bit grayscale per view Load / scaling ToTensor() $\Rightarrow\mathbf{x}_{\mathrm{raw}}\in[0,1]^{C\times 256\times 256}$ Normalization (use_norm) $\mathbf{x}_{\mathrm{norm}}=(\mathbf{x}_{\mathrm{raw}}-\mu)\oslash\sigma$ $\mu,\sigma$ estimation Randomly sample $\min(n_{\text{norm}},|{\rm train}|)$ training images without replacement (default $n_{\text{norm}}{=}512$); compute per-channel mean/std over $N{\times}H{\times}W$; clamp $\sigma\leftarrow\max(\sigma,10^{-6})$ Clipping (raw) $[0,1]$ Clipping (norm) $\big[(0-\mu)\oslash\sigma,(1-\mu)\oslash\sigma\big]$ Threat model space raw space; when attacking norm, use Eq. 13
Appendix B View Rendering and Multi-View Construction Details
This appendix details the view construction summarized in Section˜3. We follow the MODII-style design for multi-view financial image construction [11] and adopt standard time-series-to-image principles commonly used in financial prediction pipelines [4].
OHLCV rendering (ohlcv)
Each sample is rendered on a black canvas at a fixed resolution of $256\times 256$ pixels. The upper region (75% of the height) visualizes OHLC as candlestick-style glyphs (a high–low vertical stroke with open/close ticks). We shade the open–close range in gray to reflect the interval-valued view of price movements [29]. The lower region (25% of the height) encodes trading volume as a bar chart. Volumes are normalized by a rolling percentile scaler, i.e., divided by the trailing 95th percentile computed over a fixed reference window of length $L_{\mathrm{vol}}$, and then clipped to $[0,1]$.
Indicator-matrix rendering (indic)
We compute a fixed set of $N=15$ technical indicators over a lookback window of $L_{\mathrm{indic}}=15$ trading days. Concretely, we use RSI(14), WILLR(14), WMA(10), SMA(15), EMA(12), DEMA(14), TEMA(18), CCI(20), CMO(14), MACD(hist; 8,17,9), PPO(8,17; EMA), ROC(10), MFI(14), ADX(14), and SAR(0.02, 0.2); all choices and hyperparameters are fixed a priori. For each sample, we stack the most recent $L_{\mathrm{indic}}$ values into an $N\times L_{\mathrm{indic}}$ matrix (rows: indicators; columns: time) and apply per-row min–max scaling to $[0,255]$ within the current window. To keep the spatial layout consistent across splits, we compute a single indicator row order on the training split via average-linkage hierarchical clustering on the absolute Spearman correlation (distance $1-|\rho|$) with optimal leaf ordering (falling back to a spectral ordering if unavailable), and reuse it for validation/test. :contentReference[oaicite:0]index=0 :contentReference[oaicite:1]index=1 Finally, the $15\times 15$ matrix is upsampled with nearest-neighbor interpolation to $256\times 256$ (to match the image encoder input size) and rendered as a grayscale image.
Appendix C Model Architecture and Fusion Details
Table˜4 provides an implementation-level summary of all backbones and fusion heads used in the main experiments; the remainder of this appendix specifies layer-wise and fusion-head details for reproducibility.
Table 4: Implementation-level summary of model backbones and fusion heads. Here $C$ is the number of input channels (1 for single-view; 2 for early-fusion two-view). For late fusion, each branch consumes one channel and the fused head operates on concatenated features.
[…]
Notation
All models output binary logits. For single-view inputs, C=1C{=}1; for two-view inputs, C=2C{=}2 where the two views are stacked channel-wise. In late-fusion models, the input is split as $\mathbf{x}_{a}=\mathbf{x}[:,0{:}1]$ and $\mathbf{x}_{b}=\mathbf{x}[:,1{:}2]$.
Lite-CNN (PaperCNN)
Our Lite-CNN baseline is a compact CNN with three convolution blocks and two pooling operations, followed by a 2-layer MLP feature head and a final linear classifier. Concretely: (1) $\mathrm{Conv}_{3\times 3}(C\!\rightarrow\!16)$ + ReLU; (2) $\mathrm{Conv}_{3\times 3}(16\!\rightarrow\!16)$ + ReLU + $\mathrm{MaxPool}(2,2)$; (3) $\mathrm{Conv}_{3\times 3}(16\!\rightarrow\!32)$ + ReLU + $\mathrm{MaxPool}(3,3)$; then flatten and an MLP/classifier head $\mathrm{Linear}(\cdot\!\rightarrow\!256)\!\rightarrow\!\mathrm{ReLU}\!\rightarrow\!\mathrm{Dropout}\!\rightarrow\!\mathrm{Linear}(256\!\rightarrow\!128)\!\rightarrow\!\mathrm{ReLU}\!\rightarrow\!\mathrm{Linear}(128\!\rightarrow\!2)$. The flatten dimension is inferred automatically from the input resolution ($\times256$).
ResNet18-P
[…]
Late fusion with optional consistency (*-Cons)
[…]
Appendix D Additional Evaluation Details
From probabilities to labels
Unless stated otherwise, we obtain binary predictions by thresholding the predicted positive-class probability at $0.5$ and compute MCC from the resulting confusion matrix.
Reporting
[…]
Filtering protocol for min_move
When using min_move ($\tau$), we restrict evaluation to the conditional subset defined by $|r_{t+1}|\geq\tau$. Because inclusion depends on the realized next-day return, min_move is an offline benchmark definition rather than an inference-time decision rule. To keep comparisons controlled across split variants and avoid inconsistencies near split boundaries, we apply the filter once on the full timeline prior to splitting, and then run the time-block split with embargo on the filtered index set (Section˜5.4, D). The resulting sample-size impact across $\tau$ values is summarized in Table˜5.
Time-Block Split and Embargo
This paragraph complements Section˜5.4 with implementation details of the time-block split with embargo [14, 28].
Let $B$ denote the block size (block_size) and $D_{\mathrm{emb}}$ the embargo length in trading days (embargo_days). We drop $K=\lceil D_{\mathrm{emb}}/B\rceil$ adjacent blocks around split boundaries to ensure a temporal gap of at least $D_{\mathrm{emb}}$ days (possibly larger due to block granularity) [14, 28]. Table˜5 reports, for each min_move setting, the aligned sample counts before splitting and the final counts retained after applying the embargoed time-block split, together with the resulting embargo percentage. All reported results use the same split protocol across models and settings.
Table 5: Impact of min_move filtering and embargo under the time-block split protocol (block_size=30, embargo_days=20). Aligned denotes samples after filtering and view alignment, and Used denotes samples retained after applying the embargo split.
min_move Aligned Train/Val/Test Used Embargo 0 4839 3420/630/699 4749 1.86% 0.002 3666 2580/480/516 3576 2.45% 0.004 2696 1890/330/386 2606 3.34% 0.006 1989 1380/240/279 1899 4.52% 0.008 1471 1050/150/181 1381 6.12% 0.010 1054 750/90/124 964 8.54%
Appendix E Training Protocol and Reproducibility Details
All trainable baselines are implemented in PyTorch and trained under a unified optimization and checkpoint-selection protocol unless stated otherwise.
Optimization and checkpoint selection
[…]
Data loading and leakage-avoiding normalization
All images are resized to $256\times 256$ and loaded in raw $[0,1]$ space. When enabled (use_norm), we apply per-channel standardization using $(\boldsymbol{\mu},\boldsymbol{\sigma})$ estimated on the training split only; see A. Two-view inputs follow the fixed view-to-channel mapping in Section˜3.3.
Early stopping
We train for at most 40 epochs with early stopping (patience=8). We use a model-dependent minimum training epoch to stabilize optimization: min_epoch=15 for Lite-CNN and Lite-CNN-cons, and min_epoch=2 for ResNet18-P and ResNet18-P-cons.
Adversarial evaluation
Adversarial evaluation is performed at test time only under the view-aligned protocol (Section˜4.2). Attack hyperparameters and budget conversion for standardized inputs are provided in F and A.
Reproducibility
We fix random seeds for Python, NumPy, and PyTorch and log the full configuration of each run, including the selected checkpoint.
Appendix F Attack Hyperparameters and Implementation Notes
Perturbation budgets
Unless stated otherwise, perturbation budgets are defined in raw $[0,1]$ pixel space under an $\ell_{\infty}$ constraint, following standard adversarial-example evaluations [18, 1, 2]. We evaluate $\epsilon_{\mathrm{adv}}\in E$ as defined in Eq. (5), where $\epsilon_{\mathrm{adv}}=0$ corresponds to clean evaluation. When sweeping budgets, we run the evaluation once per $\epsilon_{\mathrm{adv}}$. Table˜6 summarizes the attack settings.
Loss and attack objective
All attacks are untargeted and maximize the cross-entropy loss with respect to the true label:
$\max_{\|\boldsymbol{\delta}\|_{\infty}\leq\epsilon_{\mathrm{adv}}}\ \mathcal{L}_{\mathrm{CE}}\!\left(f(\mathbf{x}+\boldsymbol{\delta}),y\right).$
Here $\mathbf{x}$ denotes the raw $[0,1]$ input; when the model consumes standardized inputs, we optimize in the standardized space using the converted budget in Eq. 13.
FGSM (one-step)
FGSM is implemented as a single-step update followed by clipping to the valid input range [1].
PGD (multi-step)
PGD performs $K=10$ iterative updates with step size $\alpha=\epsilon_{\mathrm{adv}}/K$; we use a deterministic initialization (random_start==False) from the clean input, and project after each step to enforce the $\ell_{\infty}$ constraint [2].
Table 6: Attack settings used in our experiments. Budgets are defined in raw $[0,1]$ space. When attacking standardized inputs, we convert $(\epsilon_{\mathrm{adv}},\alpha)$ using Eq. 13.
Attack Objective Budget Steps $K$ Step size $\alpha$ Random init. FGSM untargeted CE (maximize) $\epsilon_{\mathrm{adv}}\in E$ 1 $\alpha=\epsilon_{\mathrm{adv}}$ No PGD untargeted CE (maximize) $\epsilon_{\mathrm{adv}}\in E$ 10 $\alpha=\epsilon_{\mathrm{adv}}/10$ No
Appendix G Clean Baselines: Full Numerical Commentary
This appendix provides the full numerical commentary for Section˜6.1 (clean evaluation), without changing any reported values or conclusions. Representative values are summarized in Table˜1; full curves are shown in Figure˜2.
Minimum-movement filtering defines the evaluation regime
Figure˜2 shows that the minimum-movement threshold $\tau$ (Section˜5.3) largely determines whether the task is learnable on the ex-post filtered subsets defined by min_move. Without filtering or with weak filtering ($\tau\in\{0,0.002,0.004\}$), most deep models hover around zero MCC (sometimes negative), indicating near-chance behavior on the full set. As $\tau$ increases to $0.006$ and $0.008$, performance improves sharply, consistent with restricting evaluation to subsets that exclude near-zero moves and thus reduce direction-label ambiguity. Aggregating across all deep configurations (Lite-CNN/ResNet18-P and *-cons, all input modes), the mean MCC rises from 0.002-0.002 at $\tau=0$ to $0.082$ at $\tau=0.006$ and peaks at $0.092$ at $\tau=0.008$, before dropping to $0.030$ at $\tau=0.010$. The drop at $\tau=0.010$ aligns with the reduced effective sample size (Table˜5), which increases variance and destabilizes training. We stress that $\tau$ is used here to define evaluation regimes ex post, not as an inference-time filtering rule.
Lite-CNN: single-view is more robust under weak filtering; dual-view becomes beneficial once labels stabilize
Under weak filtering ($\tau=0.004$), indic already yields a clear positive signal for Lite-CNN ($0.107292\pm 0.022477$), while both early fusion is negative ($0.020004\pm 0.016396$), suggesting negative transfer when two noisy views are fused too early. Once $\tau\geq 0.006$, both becomes consistently positive and competitive (e.g., $0.088767\pm 0.042662$ at $\tau=0.006$, $0.107206\pm 0.046632$ at $\tau=0.008$), indicating that dual-view information is exploitable when labels are less noisy. At $\tau=0.006$, ohlcv is also strong ($0.109095\pm 0.039023$), implying that price-based visual patterns alone can be predictive after filtering.
Late fusion drives the main gain; consistency regularization yields modest, non-monotonic effects
Switching from early fusion (lite-cnn | both) to dual-branch late fusion without consistency ($\lambda=0$) yields a large improvement at $\tau=0.006$ (from $0.088767\pm 0.042662$ to $0.127459\pm 0.035946$), indicating that the architecture-level change (separate encoders + fusion head) is the dominant factor in our setting. Adding consistency regularization can help, but not monotonically and the gains are modest: at $\tau=0.006$, lite-cnn-cons peaks at $\lambda=0.5$ ($0.129607\pm 0.026449$), while $\lambda=1$ is lower ($0.120755\pm 0.029468$). At $\tau=0.008$, $\lambda=1$ reaches $0.103269\pm 0.031387$, close to early fusion ($0.107206\pm 0.046632$), suggesting that consistency appears to act largely as a regularizer, whose benefit is $\tau$-dependent and sensitive to the effective data regime induced by filtering.
ResNet18-P: early fusion underperforms; late fusion partially recovers but variance is higher
ResNet18-P exhibits a different regime: dual-view early fusion is consistently weak (e.g., $0.042873\pm 0.061370$ at $\tau=0.008$), while single-view modes are stronger at the same threshold (indic: $0.108023\pm 0.040145$, ohlcv: $0.109101\pm 0.072830$). Late fusion improves the dual-view setting (e.g., resnet18p-cons with $\lambda=0.5$ reaches $0.083423\pm 0.052584$ at $\tau=0.008$), but the ResNet family generally shows larger standard deviations, especially under strict filtering (e.g., resnet18p | both: $0.061769\pm 0.115898$ at $\tau=0.010$), consistent with sample-size-driven instability.
LogReg as a diagnostic baseline
LogReg can be competitive at certain thresholds (e.g., ohlcv at $\tau=0.008$ achieves $\mathrm{MCC}=0.116126$), but its behavior is highly non-monotonic across $\tau$ and input modes (e.g., both at $\tau=0.006$ yields $\mathrm{MCC}=-0.039129$). We therefore treat it primarily as a diagnostic indicator of partial linear separability after label stabilization.
Appendix H Full Branch Diagnostics for Late-Fusion Models
Branch diagnostics for late-fusion two-view models (clean test, $\tau=0.006$)
To interpret robustness trends for *-cons models, we inspect clean-test branch diagnostics that summarize cross-branch alignment (agree_ab, sym_kl with T=2T{=}2) and branch-wise versus fused MCC (Table˜7; full results in Tables 810). For Lite-CNN-Cons, enabling consistency ($\lambda>0$) sharply increases alignment (higher agree_ab, much lower sym_kl), yet branch-wise MCC remains near zero while fused-head MCC stays substantially higher. This indicates that predictive signal is expressed primarily through the feature-level late-fusion pathway rather than strong standalone branch classifiers. For ResNet18P-Cons, larger $\lambda$ likewise improves alignment and is associated with higher fused MCC, but variance remains notable and robustness failures under indicator-view attacks persist.
[…]
Model $\boldsymbol{\lambda}$ agree_ab sym_kl mcc_a / mcc_b mcc_fuse Lite-CNN-Cons 0 $0.275090\pm 0.244057$ $164898\pm 0.125751 0.0$$019377\pm 0.061519\ /\ 0.005906\pm 0.051856$ $0.127459\pm 0.035946$ Lite-CNN-Cons 0.5 $830197\pm 0.185333 0.0$$00195\pm 0.000098 0.0$$001488\pm 0.053262\ /\ -0.001753\pm 0.076580$ $0.129607\pm 0.026449$ Lite-CNN-Cons 1 $803315\pm 0.205572 0.0$$00106\pm 0.000029 0.0$$006285\pm 0.057660\ /\ -0.015335\pm 0.052686$ $0.120755\pm 0.029468$ ResNet18P-Cons 0 $439964\pm 0.247314 0.0$$74461\pm 0.065610 0.0$$14477\pm 0.047350\ /\ 0.000779\pm 0.038367 0.0$$048051\pm 0.056600$ ResNet18P-Cons 0.5 $920251\pm 0.196921 0.0$$06523\pm 0.003518 0.0$$06042\pm 0.012812\ /\ -0.002165\pm 0.015076 0.0$$067112\pm 0.031388$ ResNet18P-Cons 1 $931452\pm 0.163280 0.0$$03425\pm 0.001379 0.0$$02706\pm 0.028186\ /\ 0.009707\pm 0.019310 0.0$$095146\pm 0.057933$
Tables 810 provide confidence, class-probability means, and branch-vs-fused MCC, respectively.
[…]
Model $\boldsymbol{\lambda}$ conf_a conf_b conf_f Lite-CNN-Cons 0 $0.627352\pm 0.070257$ $0.554910\pm 0.020764$ $0.917221\pm 0.007478$ Lite-CNN-Cons 0.5 $0.508533\pm 0.004537$ $0.506999\pm 0.003842$ $0.913617\pm 0.005166$ Lite-CNN-Cons 1 $0.507229\pm 0.004763$ $0.507287\pm 0.005493$ $0.913873\pm 0.007414$ ResNet18P-Cons 0 $0.570899\pm 0.043481$ $0.567578\pm 0.040866$ $0.782159\pm 0.078004$ ResNet18P-Cons 0.5 $0.605562\pm 0.065867$ $0.611589\pm 0.066048$ $0.802972\pm 0.056561$ ResNet18P-Cons 1 $0.622989\pm 0.083976$ $0.624896\pm 0.082731$ $0.808066\pm 0.048449$
[…]
Model $\boldsymbol{\lambda}$ pos_mean_a pos_mean_b pos_mean_f Lite-CNN-Cons 0 $0.522772\pm 0.145785$ $0.511727\pm 0.058662$ $0.514203\pm 0.039776$ Lite-CNN-Cons 0.5 $0.502299\pm 0.009459$ $0.499700\pm 0.008371$ $0.511911\pm 0.056243$ Lite-CNN-Cons 1 $0.498687\pm 0.008587$ $0.498635\pm 0.009367$ $0.520245\pm 0.068696$ ResNet18P-Cons 0 $0.549286\pm 0.065936$ $0.513286\pm 0.074454$ $0.455765\pm 0.133288$ ResNet18P-Cons 0.5 $0.592447\pm 0.085277$ $0.590547\pm 0.095968$ $0.483531\pm 0.135293$ ResNet18P-Cons 1 $0.564726\pm 0.139656$ $0.564738\pm 0.140885$ $0.482701\pm 0.169748$
[…]
Model $\boldsymbol{\lambda}$ mcc_a mcc_b mcc_fuse Lite-CNN-Cons 0 $019377\pm 0.061519 0.0$$005906\pm 0.051856$ $0.127459\pm 0.035946$ Lite-CNN-Cons 0.5 $001488\pm 0.053262 0.0$$001753\pm 0.076580$ $0.129607\pm 0.026449$ Lite-CNN-Cons 1 $006285\pm 0.057660 0.0$$015335\pm 0.052686$ $0.120755\pm 0.029468$ ResNet18P-Cons 0 $14477\pm 0.047350 0.000$$0779\pm 0.038367 0.0$$048051\pm 0.056600$ ResNet18P-Cons 0.5 $006042\pm 0.012812 0.0$$02165\pm 0.015076 0.0$$067112\pm 0.031388$ ResNet18P-Cons 1 $002706\pm 0.028186 0.0$$09707\pm 0.019310 0.0$$095146\pm 0.057933$
Appendix I Sensitivity to the Operating Point: Adversarial Robustness at $\tau=0.008$
This appendix reports adversarial robustness results under a larger min_move threshold $\tau=0.008$. Because changing $\tau$ alters the labeling rule (and typically class balance), absolute clean MCC values should not be compared across $\tau$; we therefore focus on within-$\tau$ degradation patterns.
Table 11: PGD robustness summary at $\tau=0.008$ (min_move). We report MCC at two representative budgets ($0.25/255$ and $1/255$), the worst-case drop $\Delta_{\mathrm{worst}}$ over nonzero budgets $\{0.25,0.5,0.75,1\}/255$ (Eq. (8)), and the stress-test MCC at $\epsilon_{\mathrm{adv}}=2/255$.
Model Input Attack (PGD) MCC@0 MCC@0.25/255 MCC@1/255 $\boldsymbol{\Delta_{\mathrm{worst}}}$ MCC@2/255 Lite-CNN ohlcv on ohlcv 0.091 0.042 -0.101 -0.192 -0.279 Lite-CNN indic on indic 0.100 0.035 -0.187 -0.287 -0.410 Lite-CNN both (early) on ohlcv (ch0) 0.107 0.047 -0.101 -0.209 -0.300 Lite-CNN both (early) on indic (ch1) 0.107 0.043 -0.101 -0.208 -0.272 Lite-CNN both (early) on both 0.107 -0.008 -0.285 -0.393 -0.561 Lite-CNN-Cons ($\lambda{=}0$) both (late) on ohlcv (ch0) 0.100 0.061 -0.048 -0.148 -0.198 Lite-CNN-Cons ($\lambda{=}0$) both (late) on indic (ch1) 0.100 0.087 0.058 -0.043 0.032 Lite-CNN-Cons ($\lambda{=}0$) both (late) on both 0.100 0.057 -0.078 -0.178 -0.249 Lite-CNN-Cons ($\lambda{=}0.5$) both (late) on ohlcv (ch0) 0.090 0.062 -0.051 -0.141 -0.179 Lite-CNN-Cons ($\lambda{=}0.5$) both (late) on indic (ch1) 0.090 0.084 0.063 -0.027 0.043 Lite-CNN-Cons ($\lambda{=}0.5$) both (late) on both 0.090 0.053 -0.080 -0.170 -0.246 Lite-CNN-Cons ($\lambda{=}1$) both (late) on ohlcv (ch0) 0.103 0.059 -0.070 -0.173 -0.219 Lite-CNN-Cons ($\lambda{=}1$) both (late) on indic (ch1) 0.103 0.099 0.064 -0.039 0.018 Lite-CNN-Cons ($\lambda{=}1$) both (late) on both 0.103 0.042 -0.098 -0.201 -0.269 ResNet18-P ohlcv on ohlcv 0.109 -0.489 -0.977 -1.087 -0.999 ResNet18-P indic on indic 0.108 -1.000 -1.000 -1.108 -1.000 ResNet18-P both (early) on ohlcv (ch0) 0.043 -0.453 -0.962 -1.005 -1.000 ResNet18-P both (early) on indic (ch1) 0.043 -0.830 -1.000 -1.043 -1.000 ResNet18-P both (early) on both 0.043 -0.936 -1.000 -1.043 -1.000 ResNet18-P-Cons ($\lambda{=}0$) both (late) on ohlcv (ch0) 0.079 -0.240 -0.799 -0.878 -0.974 ResNet18-P-Cons ($\lambda{=}0$) both (late) on indic (ch1) 0.079 -0.927 -1.000 -1.079 -1.000 ResNet18-P-Cons ($\lambda{=}0$) both (late) on both 0.079 -0.966 -1.000 -1.079 -1.000 ResNet18-P-Cons ($\lambda{=}0.5$) both (late) on ohlcv (ch0) 0.083 -0.246 -0.791 -0.875 -0.986 ResNet18-P-Cons ($\lambda{=}0.5$) both (late) on indic (ch1) 0.083 -0.921 -1.000 -1.083 -1.000 ResNet18-P-Cons ($\lambda{=}0.5$) both (late) on both 0.083 -0.964 -1.000 -1.083 -1.000 ResNet18-P-Cons ($\lambda{=}1$) both (late) on ohlcv (ch0) 0.094 -0.247 -0.795 -0.889 -0.974 ResNet18-P-Cons ($\lambda{=}1$) both (late) on indic (ch1) 0.094 -0.944 -1.000 -1.094 -1.000 ResNet18-P-Cons ($\lambda{=}1$) both (late) on both 0.094 -0.978 -1.000 -1.094 -1.000
[…]
Model $\boldsymbol{\lambda}$ agree_ab sym_kl mcc_a / mcc_b mcc_fuse Lite-CNN-Cons 0 $0.470994\pm 0.357293$ $119371\pm 0.092575 0.0$$05352\pm 0.102980 / 0.0$$017188\pm 0.039365$ $0.100430\pm 0.034439$ Lite-CNN-Cons 0.5 $775552\pm 0.170337 0.0$$00164\pm 0.000088 0.0$$09824\pm 0.052837 / 0.0$$04969\pm 0.047469 0.0$$089973\pm 0.067209$ Lite-CNN-Cons 1 $778315\pm 0.208975 0.0$$00191\pm 0.000099 0.0$$03662\pm 0.048320 / 0.0$$007522\pm 0.051147$ $0.103269\pm 0.031387$ ResNet18P-Cons 0 $408840\pm 0.195311 0.0$$61022\pm 0.040954 0.0$$14461\pm 0.048496 / 0.0$$24219\pm 0.084118 0.0$$078607\pm 0.054662$ ResNet18P-Cons 0.5 $946823\pm 0.150407 0.0$$03367\pm 0.001974 0.0$$27162\pm 0.076826 / 0.0$$08803\pm 0.024899 0.0$$083423\pm 0.052584$ ResNet18P-Cons 1 $934392\pm 0.112509 0.0$$05748\pm 0.003453 0.0$$16075\pm 0.030005 / 0.0$$43617\pm 0.080173 0.0$$093970\pm 0.101626$
Appendix J PGD Stress Test (up to $\epsilon_{\mathrm{adv}}=2/255$)
To complement the main robustness grid (which uses $\epsilon_{\mathrm{adv}}\leq 1/255$), we extend the perturbation budget to $\epsilon_{\mathrm{adv}}=2/255$ for the late-fusion two-view Lite-CNN-Cons models. This stress test is evaluated at both $\tau=0.006$ and $\tau=0.008$. Across both operating points, Lite-CNN-Cons remains non-trivially robust under the view-constrained indic-only attack (ch1) even at $2/255$ (best mean MCC $\approx 0.048\pm 0.059$ at $\tau=0.006$ with $\lambda=1$, and $\approx 0.043\pm 0.047$ at $\tau=0.008$ with $\lambda=0.5$). In contrast, attacking ohlcv only (ch0) already drives MCC negative at the same budget (best mean MCC $\approx-0.199\pm 0.100$ at $\tau=0.006$ and $\approx-0.179\pm 0.117$ at $\tau=0.008$), and the joint two-view attack remains the hardest case (best mean MCC $\approx-0.278\pm 0.092$ at $\tau=0.006$ and $\approx-0.246\pm 0.078$ at $\tau=0.008$). These results strengthen the interpretation that late fusion primarily helps when the attacker cannot manipulate both views (redundancy across branches), while it cannot guarantee robustness under joint perturbations.
Random-start check
We also verified that enabling random_start (with multiple restarts) does not materially change the results at $2/255$ (we did not observe worse MCC), which is consistent with a locally near-linear loss landscape where PGD quickly converges to the same boundary (sign-gradient) solution regardless of initialization.
Table 13: PGD stress test at $\epsilon_{\mathrm{adv}}=2/255$. We report MCC on clean data and under three PGD threat models restricted to indic (ch1), ohlcv (ch0), or both views.
$\boldsymbol{\tau}$ $\boldsymbol{\lambda}$ MCC@0 MCC@2/255 (ch1) MCC@2/255 (ch0) MCC@2/255 (both) 0.006 0 0.127459±0.035946 0.029360±0.045407 -0.202645±0.076818 -0.285711±0.069225 0.006 0.5 0.129607±0.026449 0.030246±0.067472 -0.215792±0.097168 -0.310946±0.094411 0.006 1 0.120755±0.029468 0.048053±0.058530 -0.198704±0.099514 -0.277539±0.092451 0.008 0 0.100430±0.034439 0.031542±0.037874 -0.198313±0.070048 -0.248958±0.082638 0.008 0.5 0.089973±0.067209 0.042798±0.046734 -0.178680±0.116959 -0.246481±0.078135 0.008 1 0.103269±0.031387 0.018316±0.048887 -0.218880±0.104764 -0.268799±0.074658